Find a function whose graph is a parabola with vertex
(4, 3)
and that passes through the point
(−3, 1).

Answer by lwsshak3(11628)

Find a function whose graph is a parabola with vertex
(4, 3)and that passes through the point(−3, 1).
***
Given data shows parabola opens left.
Its basic equation: (y-k)^2=4p(x-h), (h,k)=(x,y) coordinates of vertex
(y-3)^2=4p(x-4)
(1-3)^2=4p(-3-4)
4=4p(-7)
4p=-4/7
equation:
(y-3)^2=-(4/7)(x-4)
see graph below:
y=((-4/7)(x-4))^.5+3